In​ 2008, the per capita consumption of soft drinks in Country A was reported to be 18.51 gallons. Assume that the per capita consumption of soft drinks in Country A is approximately normally​ distributed, with a mean of 18.51 gallons and a standard deviation of 5 gallons. Complete parts​ (a) through​ (d) below.

a. What is the probability that someone in Country A consumed more than 13 gallons of soft drinks in​ 2008?

​(Round to four decimal places as​ needed.)

b. What is the probability that someone in Country A consumed between 3 and 6 gallons of soft drinks in​ 2008?

​(Round to four decimal places as​ needed.)

c. What is the probability that someone in Country A consumed less than 6 gallons of soft drinks in​ 2008?

​(Round to four decimal places as​ needed.)

d. 97​% of the people in Country A consumed less than how many gallons of soft​ drinks? (2 points)

​(Round to two decimal places as​ needed.)
Answer #1

flere we are give un (18.51 52 PC X>1]) = L-P (U-18:51 CII-18:52 5 = 1-PC21 -5.5) 0.8648 698 P ( 3cnico) P (3-18:5) Cu-18.5)